• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name G.J. Lolos

    Log entry time 07:54:22 on June13,2000

    Entry number 43639

    keyword=shift summary

    Shift starts with ANOTHER trip of Q3 electron.
    This time waiting until 4:00am to go and reset.

    04:00 reset Q3 electron; seems to work fine. The voltage are ~ 30mV.
    04:40 beam is back at 62microA. Data taking started.
    05:10 the voltages for Q3 are now ~50mV
    06:02 Q3 tripped. Over temperature fault and then it quenched while the power was going down. The two voltmeters for Q3 still read "normal" readings of 58.5 and 57.7 mV.
    06:15 sent someone down to reset the magnet and it worked (cleared all faults).
    06:30 waiting for beam
    06:50 after reset, Q3 voltages went back to where they were before, and now the
    voltmeters read ~60mV
    07:10 Q3 voltages are now [60-70]mV
    07:20 beam stopped
    07:35 Q3 tripped on the "Over Temperature" fault again, and then it quenched again while the power was ramping down. Asked for an access.

    Hall A still in control access at the end of the shift.


    Time.......Run.......#evts........Current......Notes
    04:41......1997...... 1.3M ........62.5uA.......16 min run.
    04:57......1998.......1.3M.........63.5uA.......Normal running
    05:12......1999.......1.3M.........63.5uA.......Normal running
    05:27......2000.......1.3M.........63.0uA.......Two very brief beam interruptions(2min)
    05:44......2001.......1.2M.........61.8uA.......Q3 tripped. Terminated at 06:02.
    06:35......2002.......1.3M.........61.8uA........Another day in paradise. FB was OFF.
    06:51......2003.......1.2M.........62.2uA........Terminated run to turn FB on.
    07:07......2004.......0.5M.........61.9uA........FB is back ON. Beam OFF. Run over.